OpenStudy (anonymous):

PLEASE HELP! ( -3k /k-2 ) + ( 6/k-2)

OpenStudy (anonymous):

\[\frac{ -3k }{ k-2 }+\frac{ 6 }{ k-2 }\]

OpenStudy (anonymous):

What exactly do you want, simplification?

OpenStudy (anonymous):

\[\frac{ -3k }{ k-2 }+\frac{ 6 }{ k-2 } = \frac{ -3k+6 }{ k-2 } = \frac{ -3(k-2) }{ k-2 }\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Finally reduce by (k-2) and you will get -3

OpenStudy (anonymous):

Thanks !

OpenStudy (anonymous):

You are welcome